OpenStudy (anonymous):

solve using elimination method. 6x-3y=30 3x+6y=-30

OpenStudy (anonymous):

I think there needs to be another equation?

OpenStudy (anonymous):

This should be fine. You can write the first equation as: 6x = 30 + 3y 3x = 15 + 1.5y Then, substitute this in into the second equation and see what happens :)

OpenStudy (anonymous):

multiply the top equation by 2 so that you can cancel the y values.

OpenStudy (anonymous):

why by 2? thats the thing i dont understand and why the first equation? does that apply for every problem i need to solve when using the elimination method?

OpenStudy (anonymous):

yes, before he only had one equation that was why I asked that.

OpenStudy (anonymous):

@mvalerie95 u have to do that because we want to achieve the same coefficients of "6x" on the top and bottom so we can eliminate the "6x" and find the y value.

OpenStudy (anonymous):

and it does apply to every single simultaneous equation that do not have the same coefficients. see the first equation we have "6x" and the second one we only have "3x" so we have to multiply the whole bottom equation by 2, to make it to "6x"

OpenStudy (anonymous):

Basically you see a term -3y in the top equation. If you multiply the entire equation by 2 (only for this case) then you end up with a 6y term, that term is also in the second equation. It makes it easier to substitute one in the other. The more correct way would be to find an expression for y in terms of x: 6x - 3y = 30 -3y = 30 -6x y = 30/(-3) - 6x/(-3) y = -10 + 2x Then substitute this expression for y in the second equation, and you will have only x in that equation.

OpenStudy (anonymous):

so what would the sloutions be?

OpenStudy (anonymous):

solutions*

OpenStudy (anonymous):

so the new equations would be 6x-3y=30 6x+12y=-60 ?

OpenStudy (anonymous):

yes, you are right.

OpenStudy (anonymous):

I think u have understood the concept, that's good.

OpenStudy (anonymous):

and at @TimSmit, it specifically says use the "Elimination Method" so we should follow that method not by substituting.

OpenStudy (anonymous):

Now we have our two equations. 6x-3y=30 6x+12y=-60 Do you know how to do it now?

OpenStudy (anonymous):

then i add the equation which equals 12x+9y=-30?

OpenStudy (anonymous):

nope, "Elimination" we want to "Eliminate/cancel" the x.

OpenStudy (anonymous):

We will rather minus the equations.

OpenStudy (anonymous):

(6x - 6x) + (-3y-12y) = (30--60) can you do that? tell me what u get.

OpenStudy (anonymous):

u should get y = ...........

OpenStudy (anonymous):

right right subtract because we want to get rid of 6x wouldnt it be easier if i multiply by -2 and then adding the equations?

OpenStudy (anonymous):

yes sure, u can do that but minusing should be just as easy, and it seems like u struggle with minusing so u need more practice with that, try it and tell me what u get.

OpenStudy (anonymous):

6x-3y=30 -6x-12y=60 15y=60 y=4 is that correct?

OpenStudy (anonymous):

6x-3y=30 6x+12y=-60 (6x - 6x) + (-3y-12y) = (30--60)

OpenStudy (anonymous):

how would i go from there when finding the solution sets of both equations?

OpenStudy (anonymous):

have u got ur y=value?

OpenStudy (anonymous):

from there, u will figure out ur y-value of the solution to both sets. Know you have this information ( ,y) but we still need to find the x-value of the solution to both sets, so we have to substitute the y-value that u have found back into either of the original equations to solve for x. then we will have our solution to both sets (x,y)

OpenStudy (anonymous):

can you show me how to solve the whole problem step by step in your way because im following the book examples and thats why im confusing myself.

OpenStudy (anonymous):

ok back, sorry.

OpenStudy (anonymous):

here is step wise solution multiply 2nd eq. by 2 tu make the coefficients of x equal so we get 2 eq.s 6x-3y=30 6x+12y=-60 minus 2nd eq. frm 1st like -15y=90 y=-6 put in 1st eq. and you will get x =6.